If x Y z are in ap then find the value of ( x+y-Z)(-x+Y+z).
 

Grade:10

1 Answers

Khimraj
6 years ago
Let X = a-d , Y = a, Z = a+d then (X+Y-Z)(X+Y+Z) = (a-2d)(a) = {2(a-d) - a}(a) =(2X-Y)(Y) = 2XY -Y².Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.